# The Structure of Service Level Agreement of Slice-based 5G Network

### 12 Mar 2026 v2

@@ I Introduction
In legacy telecommunication networks, the requirements and characteristics of different service types are largely similar; therefore, most service level agreements (SLAs) between an operator and a tenant contain the same set of metrics. However, in slice-based 5G networks, each slice requires an individual SLA with unique elements, metrics, and structure compared to the SLAs of other slices within the same network.
As its name implies, an SLA is a formal agreement between an network operator and a tenant, or between network operators, in which the level of delivered service is precisely defined. According to the International Telecommunication Union (ITU), “the SLA is a formal agreement between two or more entities that is reached after a negotiating activity with the scope to assess service characteristics, responsibilities and priorities of every part” [4]. An SLA establishes a common understanding of a service and its relevant aspects, such as performance, availability, and responsibilities.
@@
Recently, SLAs in telecommunication networks have been extensively studied. The International Telecommunication Union (ITU) proposed a generic structure for an SLA in a multi-service-provider telecommunication environment in Recommendation E.860 [4]. The proposed SLA defines QoS-related terms and further describes the entire procedure for establishing an end-to-end SLA.
The European Telecommunications Standards Institute (ETSI) has also conducted numerous studies on SLAs, which are reported in [5], [6], and [7]. Reference [5] explores two main aspects of SLAs, namely the development phases and the SLA template, and further discusses the contents, technical features, QoS metrics and commitments, charging and billing, and reporting mechanisms of an SLA. Reference [6] investigates the SLA life cycle and penalty mechanisms, while [7] studies user demands and the various service offers provided to the tenants by a service provider.
Moreover, an end-to-end structure of QoS-oriented SLAs and a framework for real-time SLA management in multi-service packet networks are investigated in [3]. The authors presented a monitoring scheme capable of generating revenue through admission flows and calculating penalties when flows are lost. However, to the best of our knowledge, no study to date has specifically explored the SLA between a tenant and an operator in a slice-based 5G network.
In this paper, we propose an end-to-end SLA structure for a slice-based 5G communication network. We further discuss the metrics of the proposed SLA that should be considered by both the operator and the tenant during the agreement process. The rest of the paper is organized as follows. The concept and structure of the proposed SLA are thoroughly discussed in Sec. II. Subsequently, the metrics of the proposed SLA are analyzed in Sec. III. Finally, Sec. IV summarizes the main conclusions and outlines directions for future research.
@@ II The Structure of Proposed SLA
In this section, we introduce and thoroughly describe the end-to-end structure of the proposed slice-based SLA between a tenant and an operator in a 5G communication system. Moreover, we discuss two types of slice-based SLAs, namely Static SLA and Dynamic SLA, which help simplify the operation of different categories of services across various types of slices within the same 5G communication network.
@@
A major incident ($I_{\mathrm{ma}}$) requires immediate response, as it poses a serious risk to the integrity of the network, for example during traffic overload situations. A critical incident ($I_{\mathrm{cr}}$) represents the most severe condition within the slice and is typically caused by failures of hardware components.
Once an incident occurs, all monitoring metrics should be automatically examined for troubleshooting and for evaluating potential SLA breaches, as well as for determining the type of incident. If $I_{\mathrm{mi}}$ occurs within the slice, it should be resolved as soon as possible. After resolving $I_{\mathrm{mi}}$, a penalty $P$ may be calculated according to the source and severity of the incident, which the operator is required to pay to the tenant. In the case of $I_{\mathrm{mi}}$, we recommend that the operator and the tenant agree on specific threshold values for each monitoring metric associated with penalties. In this context, the tenant does not impose a penalty on the operator if the incident can be resolved without violating any of the predefined thresholds. Otherwise, the tenant imposes a penalty $P$ on the operator and formally notifies the operator to resolve the incident promptly and ensure the quality of the agreed services.
@@
Figure 1: An end-to-end structure of proposed SLA in a slicing-based network
@@ III-C3 Charge Dependent on the Time of Occurrence (Slice Level)
In some SLAs, the same outage may have different levels of importance depending on when it occurs (e.g., peak versus off-peak periods). If penalties are evaluated at the slice level (without distinguishing between subcontracts), let $\delta t$ denote the time interval between measurements, $t_{j}$ the $j$-th time instant, and $I(t_{j})$ the time-varying importance weight (dimensionless, typically $0<I(t_{j})\leq 1$). Let $T$ denote the end time of a failure period and $\Delta T$ its duration, and define $J=\Delta T/\delta t$. Then
@@
  • $J=\Delta T/\delta t$ — number of measurement intervals in the failure period.
@@ III-C4 Charge Dependent on the Time of Occurrence (Subcontracts)
When subcontracts are taken into account, a subcontract refers to an individual service component or metric commitment within the SLA (e.g., latency, throughput, or availability threshold). Let $k$ denote the number of violated subcontracts. For each subcontract $i$, let $\delta t_{i}$ denote the measurement interval, $w_{i}$ the penalty rate per unit time (at the subcontract level), $T_{i}$ the end time of a failure period, $\Delta T_{i}$ the duration of that failure period, and $I(i,t_{j})$ the importance weight at time $t_{j}$ (dimensionless, typically $0<I(i,t_{j})\leq 1$). Define $J_{i}=\Delta T_{i}/\delta t_{i}$. Then
@@
  • $J_{i}=\Delta T_{i}/\delta t_{i}$ — number of measurement intervals in the failure period of subcontract $i$.
@@ III-D Cost, Revenue, and Profit
As mentioned in Section I, each slice is designed to support a specific use case and has its own characteristics, QoS mechanisms, and architecture. Therefore, each slice must be identified by a subset of KPI requirements selected from a given set of KPIs $\mathbf{k}=[k_{1},k_{2},\ldots,k_{L}]$ through Virtual Network Function (VNF) selection.
To estimate the required volume of network resources, we consider the VNF implementation ($v$) and the size of the slice ($s$), where $s$ denotes the maximum number of user applications that can be served by the slice. Network resources may include various types, such as spectrum or bandwidth, power, time, human resources, and infrastructure. If the required amounts of these resources are represented by the vector $\mathbf{r}=[r_{1},r_{2},\ldots,r_{N}]$, where $N$ denotes the number of resource types, then by considering the cost associated with each resource, the resource requirements can be converted into the expenditure (EXP) in a manner similar to classical network cost models. Thus, we have
@@
To determine the profit generated by a slice, the cost is subtracted from the revenue, as follows:
@@
It is important to note that the KPI-to-resource mapping described in Eq. 8 is highly complex and strongly dependent on the selection of the VNF implementation ($v$). Nevertheless, since the network operator is responsible for the VNF implementation, it possesses full knowledge of this mapping. Therefore, from the operator’s perspective, it is reasonable to assume that the function $\mathbf{r}(\mathbf{k},s,v)$ is a priori known.
@@ III-E QoS-related Metrics
The definitions and measurement units of QoS-related metrics- - such as latency, delay, data rate, capacity, throughput, mobility, security, energy consumption, connection density, response time, and level of service – are predefined by standardization organizations (e.g., ITU and ETSI). As widely discussed in the literature, slice-based 5G networks aim to achieve a 1000-fold increase in system capacity, a maximum data rate of 10 Gbps and an average individual user experience of 100 Mbps, prolonged battery life through a 1000-fold reduction in energy consumption per bit, a 90% reduction in network energy usage, mobility support up to 500 km/h for high-speed users (e.g., high-speed trains), perceived availability of 99.99%, 100% coverage, and latency ranging from one millisecond to a few milliseconds [2, 1].
@@ IV Conclusions
In this paper, we presented a comprehensive end-to-end structure of an SLA between a tenant and an operator in a slice-based 5G network, aiming to balance the interests of both parties. The proposed SLA framework defines the reliability, availability, and performance of delivered telecommunication services to ensure that the right information reaches the right destination at the right time in a safe and secure manner. We also discussed key SLA metrics for slice-based networks, including throughput, penalties, cost, revenue, profit, and QoS-related metrics, which are critical factors during the agreement process. In future work, we intend to investigate different types of slice-based network SLAs, such as shared SLAs (where an SLA is shared among multiple tenants using the same slice) and hybrid SLAs (where certain tenants are prioritized while other authorized tenants may access the same slice). Moreover, this work can be extended by conducting a deeper analysis of additional QoS-related metrics, such as enhanced security mechanisms, reduced latency, and increased bandwidth.
@@
[Uncaptioned image] Bin Han (M’15 – SM’21 ) received in 2009 his B.E. degree from Shanghai Jiao Tong University, in 2012 the M.Sc. from Technische Universität Darmstadt, and in 2016 the Ph.D. (Dr.-Ing.) degree from Karlsruher Institute für Technologie. He joined Rheinland-Pfälzische Technische Universität (previously known as Technische Universität Kaiserslautern) in 2016, and is now a Senior Lecturer at its Division of Wireless Communications and Radio Navigation. Researching in the area of wireless communication and networking, he has authored over around 50 research papers and book chapters.
@@
[Uncaptioned image] Hans D. Schotten received the Diploma and Ph.D. degrees in electrical engineering from the Aachen University of Technology, Germany, in 1990 and 1997, respectively. Since August 2007, he has been a full professor and head of the Division of Wireless Communications and Radio Navigation at Technische Universität Kaiserslautern. Since 2012, he has also been Scientific Director at the German Research Center for Artificial Intelligence, heading the Intelligent Networks department. He was a senior researcher, the project manager, and the head of the research groups at Aachen University of Technology, Ericsson Corporate Research, and Qualcomm Corporate R&D. During his time at Qualcomm, he has also been the Director for Technical Standards and Coordinator of Qualcomm’s activities in European research programs.

### 13 Mar 2026 v3

@@ I Introduction
In legacy networks, the requirements of different service types are largely similar; therefore, most service level agreements (SLAs) between an operator and a tenant contain the same set of metrics. However, in 5G, each slice requires an individual SLA with unique elements, metrics, and structure compared to the SLAs of other slices within the same network.
As its name implies, an SLA is a formal agreement between a network operator and a tenant, or between network operators, in which the level of delivered service is precisely defined. According to the International Telecommunication Union (ITU), “the SLA is a formal agreement between two or more entities that is reached after a negotiating activity with the scope to assess service characteristics, responsibilities and priorities of every part” [4]. An SLA establishes a common understanding of a service and its relevant aspects, such as performance, availability, and responsibilities.
@@
Recently, SLAs in telecommunication networks have been extensively studied. The International Telecommunication Union (ITU) proposed a generic structure for an SLA in a multi-service-provider telecommunication environment in Recommendation E.860 [4]. The proposed SLA defines QoS-related terms and further describes the entire procedure for establishing an end-to-end (E2E) SLA.
The European Telecommunications Standards Institute (ETSI) has also conducted numerous studies on SLAs, which are reported in [5], [6], and [7]. Reference [5] explores two main aspects of SLAs, namely the development phases and the SLA template, and further discusses the contents, technical features, QoS metrics and commitments, charging and billing, and reporting mechanisms of an SLA. Reference [6] investigates the SLA life cycle and penalty mechanisms, while [7] studies user demands and the various services provided to the tenants by a service provider.
Moreover, an E2E structure of QoS-oriented SLAs and a framework for real-time SLA management in multi-service packet networks are investigated in [3]. The authors presented a monitoring scheme capable of generating revenue through admission flows and calculating penalties when flows are lost. However, to the best of our knowledge, no study to date has specifically explored the SLA between a tenant and an operator in a slice-based 5G network.
In this paper, we propose an E2E SLA structure for a slice-based 5G communication network. We further discuss the metrics of the proposed SLA that should be considered by both the operator and the tenant during the agreement process. The rest of the paper is organized as follows. The concept and structure of the proposed SLA are thoroughly discussed in Sec. II. Subsequently, the metrics of the proposed SLA are analyzed in Sec. III. Finally, Sec. IV summarizes the main conclusions and outlines directions for future research.
@@ II The Structure of Proposed SLA
In this section, we introduce and thoroughly describe the E2E structure of the proposed slice-based SLA between a tenant and an operator in a 5G communication system. Moreover, we discuss two types of slice-based SLAs, namely Static SLA and Dynamic SLA, which help simplify the operation of different categories of services across various types of slices within the same 5G communication network.
@@
A major incident ($I_{\mathrm{ma}}$) requires immediate response, as it poses a serious risk to the integrity of the network, for example during traffic overload situations. A critical incident ($I_{\mathrm{cr}}$) represents the most severe condition within the slice and is typically caused by failures of hardware components. Once an incident occurs, all monitoring metrics should be automatically examined for troubleshooting and for evaluating potential SLA breaches, as well as for determining the type of incident. If $I_{\mathrm{mi}}$ occurs within the slice, it should be resolved as soon as possible. After resolving $I_{\mathrm{mi}}$, a penalty $P$ may be calculated according to the source and severity of the incident, which the operator is required to pay to the tenant. In the case of $I_{\mathrm{mi}}$, we recommend that the operator and the tenant agree on specific threshold values for each monitoring metric associated with penalties. In this context, the tenant does not impose a penalty on the operator if the incident can be resolved without violating any of the predefined thresholds. Otherwise, the tenant imposes a penalty $P$ on the operator and formally notifies the operator to resolve the incident promptly and ensure the quality of the agreed services.
@@
Figure 1: An E2E structure of proposed SLA in a slicing-based network
@@ III-C3 Charge Dependent on the Time of Occurrence (Slice Level)
In some SLAs, the same outage may have different levels of importance depending on when it occurs (e.g., peak versus off-peak periods). If penalties are evaluated at the slice level (without distinguishing between subcontracts), let $\delta t$ denote the time interval between measurements, $t_{j}$ the $j$-th time instant, and $I(t_{j})$ the time-varying importance weight (dimensionless, typically $0<I(t_{j})\leq 1$). Let $T$ denote the end time of a failure period and $\Delta T$ its duration, and define $J:=\Delta T/\delta t$. Then
@@
  • $J$ — number of measurement intervals in failure period.
@@ III-C4 Charge Dependent on the Time of Occurrence (Subcontracts)
When subcontracts are taken into account, a subcontract refers to an individual service component or metric commitment within the SLA (e.g., latency, throughput, or availability threshold). Let $k$ denote the number of violated subcontracts. For each subcontract $i$, let $\delta t_{i}$ denote the measurement interval, $w_{i}$ the penalty rate per unit time (at the subcontract level), $T_{i}$ the end time of a failure period, $\Delta T_{i}$ the duration of that failure period, and $I(i,t_{j})$ the importance weight at time $t_{j}$ (dimensionless, typically $0<I(i,t_{j})\leq 1$). Define $J_{i}:=\Delta T_{i}/\delta t_{i}$. Then
@@
  • $J_{i}$ — number of measurement intervals in the failure period of subcontract $i$.
@@ III-D Cost, Revenue, and Profit
As mentioned in Section I, each slice is designed to support a specific use case and has its own characteristics, QoS mechanisms, and architecture. Therefore, each slice must be identified by a subset of KPI requirements selected from a given set of KPIs $\mathbf{k}=[k_{1},k_{2},\ldots,k_{L}]$ through network function (NF) selection.
To estimate the required volume of network resources, we consider the NF implementation ($v$) and the size of the slice ($s$), where $s$ denotes the maximum number of user applications that can be served by the slice. Network resources may include various types, such as spectrum or bandwidth, power, time, human resources, and infrastructure. If the required amounts of these resources are represented by the vector $\mathbf{r}=[r_{1},r_{2},\ldots,r_{N}]$, where $N$ denotes the number of resource types, then by considering the cost associated with each resource, the resource requirements can be converted into the expenditure (EXP) in a manner similar to classical network cost models. Thus, we have
@@
In practice, EXP(r) is computed by allocating per-slice resource usage over the SLA observation window and converting these quantities into monetary cost using operator-specific unit cost coefficients.
@@
In practice, r(k,s,v) is obtained by mapping the slice KPI requirements k, slice size s, and the chosen NF implementation v to per-slice resource allocations using the operator’s internal KPI-to-resource planning and measurement data.
@@
In practice, REV(p,s,c) is computed from the slice charging policy over the same observation window, using the agreed service price p and the measured customer demand c. To determine the profit generated by a slice, the cost is subtracted from the revenue, as follows:
@@
It is important to note that the KPI-to-resource mapping described in Eq. 8 is highly complex and strongly dependent on the selection of the NF implementation ($v$). Nevertheless, since the network operator is responsible for the NF implementation, it possesses full knowledge of this mapping. Therefore, from the operator’s perspective, it is reasonable to assume that the function $\mathbf{r}(\mathbf{k},s,v)$ is a priori known.
@@ III-E QoS-related Metrics
The definitions and measurement units of QoS-related metrics, such as latency, delay, data rate, capacity, throughput, mobility, security, energy consumption, connection density, response time, and level of service – are predefined by standardization organizations (e.g., ITU and ETSI). As widely discussed in the literature, slice-based 5G networks aim to achieve a 1000-fold increase in system capacity, a maximum data rate of 10 Gbps and an average individual user experience of 100 Mbps, prolonged battery life through a 1000-fold reduction in energy consumption per bit, a 90% reduction in network energy usage, mobility support up to 500 km/h for high-speed users (e.g., high-speed trains), perceived availability of 99.99%, 100% coverage, and latency ranging from one millisecond to a few milliseconds [2, 1].
@@ IV Conclusions
In this paper, we presented an E2E structure of an SLA between a tenant and an operator in 5G networks, aiming to balance the interests of both parties. The proposed SLA framework defines the reliability, availability, and performance of delivered telecommunication services to ensure that the right information reaches the right destination at the right time in a safe and secure manner. We also discussed key SLA metrics for slice-based networks, which are critical factors during the agreement process. In the future, we intend to investigate different types of SLAs, such as shared SLAs (where an SLA is shared among multiple tenants using the same slice) and hybrid SLAs (where certain tenants are prioritized while other authorized tenants may access the same slice). Moreover, this work can be extended by conducting a deeper analysis of additional QoS-related metrics, such as enhanced security mechanisms, reduced latency, and increased bandwidth.
@@
[Uncaptioned image] Bin Han (M’15 – SM’21 ) received in 2009 his B.E. degree from Shanghai Jiao Tong University, in 2012 the M.Sc. from Technische Universität Darmstadt, and in 2016 the Ph.D. (Dr.-Ing.) degree from Karlsruher Institute für Technologie. He joined Rheinland-Pfälzische Technische Universität (previously known as Technische Universität Kaiserslautern) in 2016, and is now a Senior Lecturer at its Division of Wireless Communications and Radio Navigation.
@@
[Uncaptioned image] Hans D. Schotten received the Diploma and Ph.D. degrees in electrical engineering from the Aachen University of Technology, Germany, in 1990 and 1997, respectively. Since August 2007, he has been a full professor and head of the Division of Wireless Communications and Radio Navigation at Technische Universität Kaiserslautern. Since 2012, he has also been Scientific Director at the German Research Center for Artificial Intelligence, heading the Intelligent Networks department. He was a senior researcher, the project manager, and the head of the research groups at Aachen University of Technology, Ericsson Corporate Research, and Qualcomm Corporate R&D.